Write two equations and graph.
Eric has already ran 5 miles and is continuing to run at a pace of 1 mile per hour. Jess has only run 2 miles and is continuing to run at a pace of 2 miles per hour. In how many hours will they have ran the same number of miles?

Found 2 solutions by stanbon, addingup:
Answer by stanbon(75887)

Write two equations and graph.
Eric has already ran 5 miles and is continuing to run at a pace of 1 mile per
hour.
Erick distance:: 5 + 1*x
------------------------------
Jess has only run 2 miles and is continuing to run at a pace of 2 miles per hour. In how many hours will they have ran the same number of miles?
Jess distance:: 2 + 2x
-------------------------------
Equation:
5 + x = 2 + 2x
Answer:: x = 3 hrs
------
Cheers,
Stan H.
-----------

Answer by addingup(3677)

5+1x = 2+2x subtract 2 and 1x on both sides:
x = 3
--------------------------
Check:
5+1(3) = 2+2(3)
5+3 = 2+6
8 = 8 Correct, in 3 hours they will have both run 8 miles